Two planes leave the Ottawa airport at the same time travelling at 920 km/h and 780 km/h. how far apart are the two planes after 3 hours, if the angle between their flight paths is 26°?


1
Expert's answer
2021-02-28T08:34:52-0500

Let distance traveled by first plane is "a", distance traveled by second plane is "b" and distance between planes is "c". If angle between flights equal to 260 using law of cosines: c2=a2+b2- 2*a*b*cos(26). So

"c=\sqrt{a^{2}+b^{2}-2*a*b*cos26}"

"a=920(km\/h)*3(h)=2760 (km)"

"b=780(km\/h)*3(h)=" "2340 (km)"

"c=\sqrt{2760^2+2340^2-2*2760*2340*cos(26)}" ,

"c\approx2176,48 (km)"